OpenStudy (anonymous):

Evaluate twenty-five minus three x for x equals zero, one, two, three, and four.

OpenStudy (anonymous):

do you need an answer for each???

OpenStudy (anonymous):

to find value of x in 25 - 3x = 0 lets add +3x to both the sides => 25 -3x + 3x = 0 + 3x => 25 = 3x now lets divide both sides with 3 => 25/3 = x ___________________________ equate 25 - 3x , equal to 1 , 2 , 3 and 4 to get the value of x in the other cases...